OpenStudy (anonymous):

Express the complex number in trigonometric form. -3i 3(cos 180° + i sin 180°) 3(cos 270° + i sin 270°) 3(cos 90° + i sin 90°) 3(cos 0° + i sin 0°)

myininaya (myininaya):

can you evaluate the following: cos(180)+i sin(180) cos(270)+i sin(270) cos(90)+ i sin(90) cos(0)+ i sin(0)

myininaya (myininaya):

which of these will result in -i

OpenStudy (anonymous):

cos(270)+i sin(270)

myininaya (myininaya):

right cos(270)=0 and sin(270)=-1 so 3(cos(270)+i sin(270))=3(0-i)=3(-i)=-3i

OpenStudy (anonymous):

ok thanks